Skip to end of metadata
Go to start of metadata

Visited LOTOS test site to swap out all three EC150s for calibration. 

HeightOriginal (Asset Panda S/N)Installed 4 Aug 21 (Asset Panda S/N)
27m1031610311
17m1030310327
7m1032210314
  • Additional Maintenance
    1. Static Dissipater at top of tower was bent 180 degrees towards the ground upon arrival. Reset to the proper position.
    2. Slight releveling
    3. Tightening of guy wires

DSM network issue continues, but Isabel is actively working on solving that.

  • No labels